# SolNet: Open-source deep learning models for photovoltaic power forecasting across the globe

### v1

@@ 5 Discussion
The findings in the result section are similar across all locations, with most of the performance gain offered by transfer learning, using the synthetic data, coming from the difference in initial performance. The learning performance is limited, and so is the asymptotic performance after a period of about six months. If data is available from a sunnier time period, this period of strong out-performance can be further shortened. This indicates that most of the difficulty is in correctly learning to forecast the seasonality inherent in PV energy production and, more importantly, the peak production. Additionally, it is important to try to mimic the situation of the target data (location, azimuth, slope) to limit the distance between source and target domain. However, the models do appear to be robust to small mis-specifications, and even exhibit improved accuracy for large mis-specifications.
5.1 Cumulative added-value
To further substantiate the significance of our research, we want to leave the reader with the following reasoning. While the asymptotic performance differential is weak between a target and transfer source model, having access to a performant model early on, when no (or limited) data is available, will have a large cumulative impact on the errors made. This can be seen in figure 8. In green we show the cumulative RMSE for the median transfer model, in blue for the median target model, and in orange for the baseline model. We can see that, over time, the impact of using the transfer model over the target model is quite large. The cumulative RMSE adds up to 1.354222Note that these RMSEs are on the zero to one scaled output for the transfer model compared to 1.954 for the target model, an improvement of over 30%. Having early access to such a model can also allow the PV owner or operator to identify commissioning or operational faults sooner.
Figure 8: The cumulative RMSE over time for different models
5.2 Limitations, and future directions
@@
In this paper, we have considered the LSTM model as the neural architecture to train the solar power forecasters. While other, more recent deep learning architectures exist (e.g. transformer models, NBeats, Nhits) that can potentially achieve even better results, the goal of this paper was to provide a principled study into transfer learning as a tool to improve performance compared to simple statistical benchmarks on the one hand while simultaneously improving sample complexity of deep neural networks. We expect the results to generalize to these other neural architectures, but this is the subject of a follow-up study, similar to how successive GPT papers grew in complexity.
@@
We have identified two key limitations of the paper. The first is the lack of access to datasets from climate zones outside of the temperate (maritime) climate zones of the Netherlands, Belgium, and even Australia, where all datasets originate from New South Wales. To further substantiate the added value of SolNet, the results shown in this study should be replicated in other parts of the world. We expect that, by making the code used for this paper freely available, more datasets will be used to validate the global performance of the transfer models. This will also allow other researchers to build on top of the models proposed here to achieve the same multiplier effect as seen in other domains such as computer vision etc.
A second limitation in the research is the lack of future covariates. The current paper exclusively used lagged PV output and cyclic variables to provide the most generalisable model possible, but adding forecasts of weather phenomena to the learning process (for example, forecasts of temperature, cloudiness or humidity) will further increase the performance of the models. Historical weather forecasts are more difficult to find however, certainly given that they have to be tailored to the specific timing and forecast horizon of the model. Together, tackling these two limitations, and integration of more deep learning models, are to be the subject of an extended, future SolNet.
@@ 6 Conclusion
This paper provides the first detailed assessment of the impact of utilising a large corpus of synthetic data in the creation of data-driven forecast models for solar power output. The improvements demonstrated hold regardless of the location of the PV panels, and are rather robust to mis-specifications in system parameters even. As evidence, the paper looked at over 300 sites in three different countries to substantiate the findings while running extensive tests lasting one year on all sites. This methodology extends existing recent work on transfer learning by minimising the need for data from ’similar sites’, and shows that there is a large boost to the initial predictive performance. As expected, this differential subsides over time, and is almost zero after a one year period, but the impact of using transfer models is accumulative over time, giving users a powerful tool to improve their forecasting performance at the early onset of (or even before) installing a new PV setup. The code, along with a functioning example for using SolNet for solar forecasting anywhere on land around the globe, has been open-sourced [^3] https://anonymous.4open.science/r/SolNet-CD07/README.md. We expect this work to be the first of many that have the potential to lead to general-purpose deep learning models for energy forecasting.

### v2

@@ 5 Discussion
The findings in the result section are similar across all locations, with most of the performance gain offered by transfer learning, using the synthetic data, coming from the difference in initial performance. This provides for a useful tool for any PV installation with limited to no data available. Using only open-source historical PV and weather data, and ideally some meta-data on the actual PV installations, a model can be made to give accurate zero-shot PV predictions anywhere in the world.
@@
There are several important considerations to be made to assess the value of transfer learning. Weather data has a strong positive effect on the predictive performance of models, and the translation from actuals in the source model as a proxy for forecasts, to actual forecasts holds up. If no data is available from a sunnier time period, the transfer model will have considerably more added value that will last longer, although here as well, weather data in the transfer model mitigates this effect. This indicates that most of the difficulty is in correctly learning to forecast the seasonality inherent in PV energy production and, more importantly, the peak production. Finally, it is important to try to mimic the situation of the target data (location, azimuth, slope) to limit the distance between source and target domain. However, the transfer models do appear to be robust to small mis-specifications, but larger mis-specifications can have a detrimental effect, especially when detailed external covariates, tied with the location, are used for predictive purposes.
5.1 Future directions and limitations
@@
In this paper, we have considered the LSTM model as the neural architecture to train the solar power forecasters. While other, more recent deep learning architectures exist (e.g. transformer models, NBeats, Nhits) that can potentially achieve even better results, the goal of this paper was to provide a principled study into transfer learning as a tool to improve performance compared to simple statistical benchmarks on the one hand while simultaneously improving sample complexity of deep neural networks. We expect the results to generalize to these other neural architectures, but this is the subject of a follow-up study.
@@
One limitation of the paper is the lack of access to datasets from climate zones outside of the temperate (maritime) climate zones of the Netherlands, Belgium, and even Australia, where all datasets originate from New South Wales. To further substantiate the added value of SolNet, the results shown in this study should be replicated in other parts of the world. We expect that, by making the code used for this paper freely available, more datasets will be used to validate the global performance of transfer models. This will also allow other researchers to build on top of the models proposed here to achieve the same multiplier effect as seen in other domains such as computer vision etc.
@@ 6 Conclusion
This paper provides a detailed assessment of the impact of utilising a large corpus of open-source (synthetic) data in the creation of data-driven forecast models for solar power output. The improvements demonstrated hold regardless of the location of the PV panels, and are rather robust to mis-specifications in system parameters. As evidence, the paper looked at over 300 sites in three different countries to substantiate the findings while running extensive tests lasting one year on all sites.
This methodology extends existing recent work on transfer learning by minimising the need for data from ’similar sites’, and shows that there is a large boost to the initial predictive performance. As expected, this differential subsides over time, and is almost zero after a one year period, but the impact of using transfer models is accumulative over time and dependent on several other factors like seasonality, giving users a powerful tool and a set of guidelines to improve their forecasting performance at the early onset of (or even before) installing a new PV setup.
The code, along with a functioning example for using SolNet for solar forecasting anywhere on land around the globe, has been open-sourced [^3] https://anonymous.4open.science/r/SolNet-CD07/README.md. We expect this work to be the first of many that have the potential to lead to general-purpose deep learning models for energy forecasting.